# Reloc3r: Large-Scale Training of Relative Camera Pose Regression for Generalizable, Fast, and Accurate Visual Localization

### v1

@@ Abstract
Visual localization aims to determine the camera pose of a query image relative to a database of posed images. In recent years, deep neural networks that directly regress camera poses have gained popularity due to their fast inference capabilities. However, existing methods struggle to either generalize well to new scenes or provide accurate camera pose estimates. To address these issues, we present Reloc3r, a simple yet effective visual localization framework. It consists of an elegantly designed relative pose regression network, and a minimalist motion averaging module for absolute pose estimation. Trained on approximately 8 million posed image pairs, Reloc3r achieves surprisingly good performance and generalization ability. We conduct extensive experiments on 6 public datasets, consistently demonstrating the effectiveness and efficiency of the proposed method. It provides high-quality camera pose estimates in real time and generalizes to novel scenes. Code, weights, and data at: https://github.com/ffrivera0/reloc3r.
@@ 1 Introduction
Absolute pose regression (APR) approaches [42, 101, 43, 69, 114, 65, 18, 102, 88, 121, 89] directly regress camera poses from images, offering much faster inference times and high accuracy. However, these methods are inherently scene-specific and typically require dense viewpoint coverage during training, limiting their real-world applicability. Recent attempts [68, 20, 21, 57, 22] to enhance accuracy through synthetic data generation introduce significant computational overhead, hampering widespread deployment. In contrast, relative pose regression (RPR) approaches [48, 7, 35, 1, 4, 48, 126, 113] estimate relative poses between database-query image pairs. These methods mitigate the need for per-scene training while maintaining the test-time efficiency of APR models. However, even the most advanced RPR methods [29, 99, 113, 4] have not yet matched APR approaches in localization accuracy. While several RPR methods [48, 99, 113, 4] have shown the ability to generalize across different datasets, this often results in further reduced camera pose accuracy. Therefore, most of the aforementioned approaches struggle with one of the following three criteria: novel-scene generalization, test-time efficiency, and camera pose accuracy.
@@
  • Comprehensive experiments across 6 popular evaluation datasets consistently demonstrate the effectiveness of our proposed approach.
@@ 2 Related Work
RPR approaches aim to generalize across different scenes by learning the relative pose between image pairs. Localization is achieved by regressing the relative pose between the query image and the most similar (or top-$K$) database images. The metric scale of the relative translation can be estimated approximately from a single database-query pair [7, 35, 1, 4], while more precise absolute positioning is possible through multi-view triangulation [48, 126, 113, 31]. Despite this capability, the best RPR methods [29, 126, 113, 4] lag significantly behind APR methods. Furthermore, the generalization ability of existing RPR models remains limited. For instance, methods like Relative PN [48] and Relpose-GNN [99] can adapt to new datasets and scenes, but their localization errors nearly double in these scenarios. Map-free [4], despite being trained on large datasets (approximately 523k samples), still relies on separate models for indoor and outdoor settings. Other research [115, 124, 104, 56, 125] has explored joint multi-view pose estimation through pose regression. However, these approaches remain dependent on dataset-specific training, which limits their scalability.
@@ 3 Method
Method overview. Figure 2 illustrates an overview of the proposed visual localization method. It comprises two main components: a relative pose regression network, and a motion averaging module. The query image $I_{q}$ is paired with the top-$K$ database images using an off-the-shelf image retrieval approach [3], creating a set of image pairs $\textbf{Q}=\{(I_{d_{k}},I_{q})\ |\ k=1,...,K\}$. The relative pose regression network processes each image pair from Q independently to determine their relative poses $P_{d,q}$ and $P_{q,d}$. With known poses of database images $\{\hat{P}{d_{1}},...,\hat{P}{d_{K}}\}$, it is possible to infer the query image’s absolute rotation from a single pair. However, these estimates are noisy, and the translation vector’s metric scale remains uncertain. To tackle these issues, a motion averaging module performs rotation averaging and camera center triangulation, eventually producing the absolute metric pose of the query image.
@@ 3.1 Relative Camera Pose Regression
The proposed relative camera pose regression network, inspired by DUSt3R [109], takes an image pair $I_{1},I_{2}$ as input. While we assume the images have the same resolution (consistent with DUSt3R), their resolutions may differ in practice. The network divides the images into patches and processes them as tokens through a Vision Transformer (ViT) encoder. A ViT decoder then uses a cross-attention mechanism to exchange information between tokens from both branches. This is followed by regression heads that predict the relative poses $\hat{P}_{I_{1},I_{2}}$ and $\hat{P}_{I_{2},I_{1}}$. Notably, the two branches are fully symmetrical with shared weights.
@@ 4 Experiments
Training data. To build up-to-scale training pairs with ground-truth relative poses, similar to DUSt3R [109], we process $\sim$8 million image pairs from 7 public datasets. These datasets cover a range of scenes, from object-centric to indoor and outdoor environments. We present the statistics of our training data in Table 1. We convert the ground-truth relative poses to a unified format: relative coordinate transformation aligned with the OpenCV [17] definition. Each image is center-cropped based on its principal points and resized to a width of 512 pixels.
@@
Table 1: Training data of Reloc3r. It comprises about 8 million image pairs from 7 public sources, covering a wide range of scenes from object-centric to indoor and outdoor environments.
Implementation details. By default, the proposed Reloc3r uses $m=24$ encoder blocks, $n=12$ decoder blocks, followed by the pose regression head with $h=2$ convolutional layers. To improve memory usage and speed in self and cross-attention, we employ memory-efficient attention. This design achieves approximately a 14% speed increase while saving 25% of GPU memory during training. We initialize Reloc3r with DUSt3R’s pre-trained 512-DPT weights. For decoder initialization, we use the weights from DUSt3R decoder2, as it’s pre-trained to perform coordinate transformation. The full model is trained on 8 AMD MI250x-40G GPUs with a batch size of 8 and a learning rate starting at 1e-5, decaying to 1e-7. More details are presented in the supplementary material.
For the visual localization task, following the literature [77, 99], we apply NetVLAD [3] for image retrieval and use the top 10 similar image pairs. We directly use these retrieved image pairs without distance-based clustering, filtering, or other heuristics to enhance their spatial distribution. All evaluations are conducted on a 24GB NVIDIA GeForce RTX 4090 GPU.
@@ 4.1 Relative Camera Pose Estimation
Reloc3r-512 (Ours) 34.79 58.37 75.56 66.70 80.20 88.39 38.18 56.39 70.34 42 ms
Table 3: Relative camera pose evaluation on the ScanNet1500 [26, 78], RealEstate10K [127] and ACID [61] datasets. The best results for each method category are highlighted in bold. Our method outperforms all the pose regression competitors. Moreover, it achieves the best results (underlined) among all the competitors on several datasets and metrics. Notably, our method runs in real-time, which is over 50×50$\times$50 × faster than the state-of-the-art Non-PR method.
@@
ImageNet+NCM [126$\dagger$ - - - - - - - 0.19 / 4.30 None
@@
ImageNet+NCM [126$\dagger$ - - - - - 0.83 / 1.36 - None
@@
Since no existing pose regression (PR) methods have been specifically designed for these datasets, we primarily compare our approach with non-PR methods. Additionally, we evaluate the current state-of-the-art relative pose regression methods, ExReNet [113] and Map-free [4], on these datasets. The results are presented in Table 3. Reloc3r outperforms other PR methods by a significant margin across all three datasets. When compared to non-PR methods, Reloc3r also achieves SoTA performance. Specifically, on the ScanNet1500 dataset, Reloc3r delivers top performance with AUC at all thresholds, outperforming its baseline DUSt3R [109] by around 13% on AUC@20. On the RealEstate10K dataset, Reloc3r achieves performance comparable to NoPoSplat [119], with AUC values of 66.70 / 80.20 / 88.39% at thresholds of 5, 10, and 20, respectively. Similarly, on the ACID dataset, Reloc3r achieves high accuracy (AUC@20 = 70.34%), demonstrating its generalization ability to unseen datasets. Furthermore, Reloc3r runs at an inference time of just 42 ms at an image resolution of 512 in width. This is significantly faster than many non-PR methods such as NoPoSplat [119] ($>$2000 ms) and ROMA [34] (300 ms), and on par with PR methods. Such speed makes Reloc3r ideal for real-time applications.
@@ 4.3 Analyses
Table 6: Ablation studies examining asymmetric network architecture and pose prediction with metric scale.
Figure 4: We present interesting findings: the cross-attention layers in Reloc3r appear to match patch correspondences between image pairs. The top row showcases matches from Efficient LoFTR, while the bottom row displays the top-3 cross-attention responses from Reloc3r’s decoder. We observe that the correlated regions in Reloc3r are superior to those of Efficient LoFTR, even though Reloc3r is trained solely with pose supervision.
@@
Comparison with pose prediction with metric scale. We train a version of Reloc3r-512 with metric poses as output. The lower accuracy results in Table 6 validate the effectiveness of our default non-metric design.
Interesting findings. We seek to understand why Reloc3r achieves accurate pose estimation. Drawing inspiration from feature-matching methods, we visualize several cross-attention maps in Reloc3r’s decoders. Interestingly, we discover that some cross-attention layers have developed the ability to match patch correspondences. Figure 4 illustrates an example of this phenomenon.
Limitations. A failure case of Reloc3r occurs when the query image and all the retrieved database images are perfectly collinear, leading to a degeneracy issue. In this scenario, the metric scale becomes unsolvable using any motion averaging methods. Fortunately, this situation rarely occurs in our evaluations.
Please refer to the supplementary material for more details and discussions of the experiments above.
@@ 5 Conclusion
In this paper, we present Reloc3r, a simple yet effective visual localization framework. It consists of an elegantly designed relative pose regression network with a minimalist motion averaging module. Leveraging large-scale training on around 8 million image pairs, Reloc3r demonstrates strong generalization capability, high efficiency, and accurate pose estimation performance across multiple datasets, while remaining parsimonious. At a minimum, it sets a strong baseline for pose regression-based visual localization. Future research could further explore the scale and diversity of data, as well as different network architectures.
@@ Appendix B Detailed Ablation Studies
The asymmetric version of Reloc3r follows DUSt3R’s design [109], which employs separate decoders and regression heads for the two input images. However, this approach increases the number of learnable parameters and introduces a potential bias based on the image order. To mitigate this bias, DUSt3R incorporates flipped image pairs during training, which adds additional computational overhead. As shown in Table 6 in the main paper, we demonstrate that the asymmetric version performs even worse than the default Reloc3r on the ScanNet1500 dataset [26, 78]. This underscores the benefits of our fully symmetric architecture, where both branches share decoder and prediction head. Remarkably, our model (with 0.42B parameters) achieves superior accuracy while using approximately 28% fewer parameters compared to the asymmetric variant.
@@ Appendix C More Analyses
Table 8: Additional results on the Cambridge Landmarks [44].
@@ Appendix D Additional Comparisons
In our evaluation of metric pose estimation, we compare results with and without motion averaging. Due to the challenge of learning metric scales, using top-2 motion averaging yields significantly better results compared to single pairs. For Reloc3r-512, we test varying numbers of top-$K$ image pairs. While increasing the number of images reduces error, it also leads to longer inference times. We also try to adopt LazyLoc [31]’s rotation and translation averaging modules as robust estimators. These provide limited improvements across most scenes, with the notable exception of GreatCourt, which features extensive repetitive patterns and similar regions. Since Reloc3r does not produce matches, it cannot adopt the post-optimization step used in LazyLoc. Like other pose regression-based methods, Reloc3r therefore still underperforms in pose accuracy compared to SoTA feature matching-based methods on large-scale scenes.

### v2

@@ Abstract
Visual localization aims to determine the camera pose of a query image relative to a database of posed images. In recent years, deep neural networks that directly regress camera poses have gained popularity due to their fast inference capabilities. However, existing methods struggle to either generalize well to new scenes or provide accurate camera pose estimates. To address these issues, we present Reloc3r, a simple yet effective visual localization framework. It consists of an elegantly designed relative pose regression network, and a minimalist motion averaging module for absolute pose estimation. Trained on approximately eight million posed image pairs, Reloc3r achieves surprisingly good performance and generalization ability. We conduct extensive experiments on six public datasets, consistently demonstrating the effectiveness and efficiency of the proposed method. It provides high-quality camera pose estimates in real time and generalizes to novel scenes. Code: https://github.com/ffrivera0/reloc3r.
@@ 1 Introduction
Absolute pose regression (APR) approaches [42, 103, 43, 70, 117, 66, 18, 104, 89, 124, 90] directly regress camera poses from images, offering much faster inference times and high accuracy. However, most of these methods are inherently scene-specific and typically require dense viewpoint coverage during training, limiting their real-world applicability. Recent attempts [69, 20, 21, 58, 22] to enhance accuracy through synthetic data generation introduce significant computational overhead, hampering widespread deployment. In contrast, relative pose regression (RPR) approaches [49, 7, 35, 1, 4, 49, 129, 116] estimate relative poses between database-query image pairs. These methods mitigate the need for per-scene training while maintaining the test-time efficiency of APR models. However, even the most advanced RPR methods [29, 101, 116, 4] have not yet matched APR approaches in localization accuracy. While several RPR methods [49, 101, 116, 4] have shown the ability to generalize across different datasets, this often results in further reduced camera pose accuracy. Therefore, most of the aforementioned approaches struggle with one of the following three criteria: novel-scene generalization, test-time efficiency, and camera pose accuracy.
@@
  • Comprehensive experiments across six popular evaluation datasets consistently demonstrate the effectiveness of our proposed approach.
@@ 2 Related Work
RPR approaches aim to generalize across different scenes by learning the relative pose between image pairs. Localization is achieved by regressing the relative pose between the query image and the most similar (or top-$K$) database images. The metric scale of the relative translation can be estimated approximately from a single database-query pair [7, 35, 1, 4], while more precise absolute positioning is possible through multi-view triangulation [49, 129, 116, 31]. Despite this capability, the best RPR methods [29, 129, 116, 4, 45] lag significantly behind APR methods. Furthermore, the generalization ability of existing RPR models remains limited. For instance, methods like Relative PN [49] and Relpose-GNN [101] can adapt to new datasets and scenes, but their localization errors nearly double in these scenarios. Map-free [4], despite being trained on large datasets (about 523K samples), still relies on separate models for indoor and outdoor settings. Other approaches explore multi-view pose estimation [118, 127, 106, 57, 128], as well as wide-baseline [77] and panoramic methods [100]. However, they remain dataset-specific training, which limits their scalability.
@@ 3 Method
Method overview. Figure 2 illustrates an overview of the proposed visual localization method. It comprises two main components: a relative pose regression network, and a motion averaging module. The query image $I_{q}$ is paired with the top-$K$ database images using an off-the-shelf image retrieval approach [3], creating a set of image pairs $\textbf{Q}=\{(I_{d_{k}},I_{q})\ |\ k=1,...,K\}$. The relative pose regression network processes each image pair from Q independently to determine their relative poses $P_{d,q}$ and $P_{q,d}$. With known poses of database images $\{\hat{P}{d_{1}},...,\hat{P}{d_{K}}\}$, it is possible to infer the query image’s absolute rotation from a single pair. However, these estimates can be noisy, and the translation vector’s metric scale remains uncertain. To tackle these issues, a motion averaging module performs rotation averaging and camera center triangulation, eventually producing the absolute metric pose of the query image.
@@ 3.1 Relative Camera Pose Regression
The proposed relative camera pose regression network, inspired by DUSt3R [111], takes an image pair $I_{1},I_{2}$ as input. While we assume the images have the same resolution (consistent with DUSt3R), their resolutions may differ in practice. The network divides the images into patches and processes them as tokens through a Vision Transformer (ViT) encoder. A ViT decoder [113, 114] then uses a cross-attention mechanism to exchange information between tokens from both branches. This is followed by regression heads that predict the relative poses $\hat{P}_{I_{1},I_{2}}$ and $\hat{P}_{I_{2},I_{1}}$. The two branches are fully symmetrical with shared weights.
@@ 4 Experiments
Training data. To build up-to-scale training pairs with ground-truth relative poses, similar to DUSt3R [111], we process around eight million image pairs from seven public datasets. The statistics are reported in Table 1. We convert the ground-truth relative poses to a unified format: relative coordinate transformation aligned with the OpenCV [17] definition. Each image is center-cropped based on its principal points and resized to a width of 512 pixels.
@@
Table 1: Training data of Reloc3r. It comprises about eight million image pairs from seven public sources, covering a wide range of scenes from object-centric to indoor and outdoor environments.
Implementation details. By default, the proposed Reloc3r uses $m=24$ encoder blocks, $n=12$ decoder blocks, followed by the pose regression head with $h=2$ convolutional layers. To improve memory usage and speed in self and cross-attention, we employ memory-efficient attention. This design achieves approximately a 14% speed increase while saving 25% of GPU memory during training. We initialize Reloc3r with DUSt3R’s pre-trained 512-DPT weights. For decoder initialization, we use the weights from DUSt3R decoder2, as it’s pre-trained to perform coordinate transformation. The full model is trained on 8 AMD MI250x-40G GPUs with a batch size of 8 and a learning rate starting at 1e-5, decaying to 1e-7.
For the visual localization task, following the literature [79, 101], we apply NetVLAD [3] for image retrieval and use the top 10 similar image pairs. We directly use these retrieved image pairs without distance-based clustering, filtering, or other heuristics to enhance their spatial distribution. All evaluations are conducted on a 24GB NVIDIA GeForce RTX 4090 GPU and we use mixed precision fp16 / fp32 to improve speed and memory efficiency without sacrificing accuracy.
@@ 4.1 Relative Camera Pose Estimation
Reloc3r-512 (Ours) 34.79 58.37 75.56 66.70 80.20 88.39 38.18 56.39 70.34 25 ms
Table 3: Relative camera pose evaluation on the ScanNet1500 [26, 80], RealEstate10K [130] and ACID [62] datasets. The best results for each method category are highlighted in bold. Our method outperforms all the pose regression competitors. Moreover, it achieves the best results (underlined) among all the competitors on several datasets and metrics. Notably, our method runs in real-time, which is over 50×50$\times$50 × faster than the state-of-the-art Non-PR method. Please refer to Figure 1 for an intuitive comparison with representative methods.
@@
ImageNet+NCM [129$\dagger$ - - - - - - - 0.19 / 4.30 None
@@
ImageNet+NCM [129$\dagger$ - - - - - 0.83 / 1.36 - None
@@
Since no existing pose regression (PR) methods have been specifically designed for these datasets, we primarily compare our approach with non-PR methods. Additionally, we evaluate the current state-of-the-art relative pose regression methods, ExReNet [116, 93] and Map-free [4], on these datasets. The results are presented in Table 3. Reloc3r outperforms other PR methods by a significant margin across all three datasets. When compared to non-PR methods, Reloc3r also achieves SoTA performance. Specifically, on the ScanNet1500 dataset, Reloc3r delivers top performance with AUC at all thresholds, outperforming its baseline DUSt3R [111] by around 13% on AUC@20. On the RealEstate10K dataset, Reloc3r achieves performance comparable to NoPoSplat [122], with AUC values of 66.70 / 80.20 / 88.39% at thresholds of 5, 10, and 20, respectively. Similarly, on the ACID dataset, Reloc3r achieves high accuracy (AUC@20 = 70.34%), demonstrating its generalization ability to unseen datasets. Furthermore, Reloc3r runs at an inference time of just 42 ms at an image resolution of 512 in width. This is significantly faster than many non-PR methods such as NoPoSplat [122] ($>$2000 ms) and ROMA [34] (300 ms), and on par with PR methods. Such speed makes Reloc3r ideal for real-time applications.
@@ 4.3 Analyses
Table 6: Ablation studies examining asymmetric network architecture and pose prediction with metric scales.
Figure 4: The top row showcases matches from Efficient LoFTR, while the bottom row displays the top-3 cross-attention responses from Reloc3r’s decoder. We observe that the correlated regions in Reloc3r are superior to those of Efficient LoFTR, even though Reloc3r is trained solely with pose supervision.
We summarize most insights here and refer the reader to the supplementary material for more details and discussions.
@@
Comparison with pose prediction with metric scale. We train a version of Reloc3r-512 with metric poses as output. The lower accuracy results reported in Table 6 validate the effectiveness of our default design without metric scales.
Interesting findings. We visualize the cross-attention responses in some blocks from Reloc3r’s decoder. Interestingly, we find that several layers have developed the ability to match patch correspondences, as illustrated in Figure 4.
Limitations. A failure case of Reloc3r occurs when the query image and all the retrieved database images are perfectly collinear, leading to a degeneracy issue that makes the metric scale unsolvable using motion averaging methods.
@@ 5 Conclusion
In this paper, we present Reloc3r, a simple yet effective visual localization framework. It consists of an elegantly designed relative pose regression network with a minimalist motion averaging module. Leveraging large-scale training on around eight million image pairs, Reloc3r demonstrates strong generalization capability, high efficiency, and accurate pose estimation performance across multiple datasets, while remaining parsimonious. We hope it will advance research on data scale and diversity in visual localization.
Acknowledgements. This work is supported by the Early Career Scheme of the Research Grants Council (grant # 27207224), the HKU-100 Award, a donation from the Musketeers Foundation, and in part by the JC STEM Lab of Autonomous Intelligent Systems funded by The Hong Kong Jockey Club Charities Trust. This work is also supported in part by the Research Council of Finland for funding the projects (grants No. 353138, No. 362407, and No. 352788). We acknowledge the computational resources provided by the CSC-IT Center for Science, Finland and the HKU-Shanghai ICRC. We thank the reviewers for their valuable feedback. Siyan Dong would like to thank the support from HKU School of Computing and Data Science.
@@ Appendix B Detailed Ablation Studies
The asymmetric version of Reloc3r follows DUSt3R’s design [111], which employs separate decoders and regression heads for the two input images. However, this approach increases the number of learnable parameters and introduces a potential bias based on the image order. To mitigate this bias, DUSt3R incorporates flipped image pairs during training, which adds additional computational overhead. As shown in Table 6 in the main paper, we demonstrate that the asymmetric version performs even worse than the default Reloc3r on the ScanNet1500 dataset [26, 80]. This underscores the benefits of our fully symmetric architecture, where both branches share decoder and prediction head. Remarkably, our model (with 0.43B parameters) achieves superior accuracy while using approximately 28% fewer parameters compared to the asymmetric variant.
@@
Rotation representations. We use a continuous 9D-to-SO(3) mapping [52] in Reloc3r to avoid the discontinuities found in 3D and 4D representations. In Table 7, we reports an ablation study using different rotation representations. The experiments are trained on ScanNet++ [123] and tested on ScanNet1500 [26, 80]. The results demonstrate the effectiveness of the 9D rotation representation.
Rot. representations 3D 4D 9D (default)
AUC@20 66.81 67.87 68.70
Table 7: Ablation study for different rotation representations.
@@ Appendix C More Analyses
Table 9: Additional results on the Cambridge Landmarks [44]. Note that although DUSt3R-512 regresses coordinates, it performs pixel-to-pixel matching with these regressed coordinates for accurate visual localization. The inference times of Reloc3r are reported using fp32.
@@
Model sizes. Previous works mainly focus on algorithm design, yet we take a different direction by scaling up the training to develop (to the best of our knowledge) the first foundation model for camera pose regression. As a result, Reloc3r’s relative pose regression network contains 0.43B parameters - far larger than existing camera pose regression networks (e.g., Map-free with 22M and Marepo with 10M parameters). Despite its size, it achieves real-time inference on consumer-grade GPUs like NVIDIA 3090/4090. We chose Transformer architectures as our backbone for their proven ability to scale better than Convolutional Neural Networks (CNNs). Our experiments with Map-free (ResUNet) showed that its 22M parameters led to underfitting on our training data. Even after expanding the CNN’s Resblocks and feature dimensions (up to 0.1B parameters), the model only memorized the training data. All CNN models we tested performed poorly, achieving AUC@20 $<$5 on the ScanNet1500 datasetet. While their rotation accuracy can be reasonable, their translation accuracy is poor.
Scale and diversity of training data. In Table 10, we show that larger training sets consistently improve pose estimation accuracy. Removing domain-specific data (such as the object-centric Co3Dv2 dataset) has minimal impact on accuracy in other domains. This suggests that diverse data helps with generalization, while domain-specific data improves accuracy within its domain.
AUC@20 on datasets ScanNet1500 RE10K ACID
Reloc3r-512 trained w/ ScanNet++ only 68.70 58.52 51.15
Reloc3r-512 trained w/o RE10K & Co3Dv2 75.46 84.44 67.41
Reloc3r-512 trained w/o RE10K 75.55 85.33 67.76
Reloc3r-512 full training 75.56 88.39 70.34
Table 10: Ablation study on training data.
@@ Appendix D Additional Comparisons
Comparison with FAR [77]. Recent works FAR [77] and PanoPose [100] design pose regression networks for wide baseline pairs and panorama images. While FAR performs well on images with few overlaps, it underperforms Reloc3r on popular datasets used in the main paper. Specifically, we tested FAR on ScanNet1500, RE10K, and ACID datasets, achieving AUC@20 of 28.19, 37.67, and 44.98%, respectively. Since PanoPose has not released its code yet, we look forward to comparing with it in the future.
@@
In our evaluation of metric pose estimation, we compare results with and without motion averaging. Due to the challenge of learning metric scales, using top-2 motion averaging yields significantly better results compared to single pairs. For Reloc3r-512, we test varying numbers of top-$K$ image pairs. While increasing the number of images reduces error, it also leads to longer inference times. We also try to adopt LazyLoc [31]’s rotation and translation averaging modules as robust estimators. These provide limited improvements across most scenes, with the notable exception of GreatCourt, which features extensive repetitive patterns and similar regions. Since Reloc3r does not produce matches, it cannot adopt the post-optimization step used in LazyLoc. Like other pose regression-based methods, Reloc3r therefore still underperforms in pose accuracy compared to SoTA feature matching-based methods on large-scale scenes. The accuracy of pose regression also can not match with those of scene coordinate regression (SCR) based methods, as SCR methods typically require per-scene training and can take long inference times.